# Well-balanced high order finite difference WENO schemes for a first-order Z4 formulation of the Einstein field equations

### v1

@@ 4.4 Stationary black holes
Finally, the bottom right panels of all figures from Fig. 10 to Fig. 13 monitor the evolution of the Einstein constraints, which remain perfectly constant all along the simulation. This is a remarkable result that can only be obtained thanks to the well balancing of the numerical scheme.
Figure 10: Schwarzschild black hole: Figs. 10a-10d show $\alpha$, $K_{zz}$ and $\gamma_{zz}$ profiles for the Schwarzschild black hole ($a=0$). Fig. 10e shows the evolution of ADM-constraints in terms of the $L_{2}$-norm. The final time is $t=1000$, and we have used a seventh-order accurate FD-WENO scheme.
Figure 11: Schwarzschild black hole: Figs. 11a-11d show $\alpha$, $K_{zz}$ and $\gamma_{zz}$ profiles for the Schwarzschild black hole ($a=0$). Fig. 11e shows the evolution of ADM-constraints in terms of the $L_{2}$-norm. The final time is $t=1000$, and we have used a sixth-order accurate AFD-WENO scheme.
Figure 12: Kerr black hole: Figs. 12a-12d show $\alpha$, $K_{zz}$ and $\gamma_{zz}$ profiles for the Kerr black hole ($a=0.9$). Fig. 12e shows the evolution of ADM-constraints in terms of the $L_{2}$-norm. The final time is $t=1000$, and we have used a seventh-order accurate FD-WENO scheme.
Figure 13: Kerr black hole: Figs. 13a-13d show $\alpha$, $K_{zz}$ and $\gamma_{zz}$ profiles for the Kerr black hole ($a=0.9$). Fig. 13e shows the evolution of ADM-constraints in terms of the $L_{2}$-norm. The final time is $t=1000$, and we have used a sixth-order accurate AFD-WENO scheme.
@@ 4.5 Head on collision of two black holes
The correct simulation of binary mergers (either black holes or neutron stars) is of course the final goal of numerical relativity, which is proving to be an invaluable scientific tool in the preset gravitational wave era (Abbott et al., 2019). In this work, still at the level of an exploratory calculation, before a dedicated analysis is performed, we have considered the head-on collision of two nonrotating black holes, which are modelled as two moving punctures. We compute the initial conditions through the TwoPunctures initial data code of Ansorg et al. (2004), selecting two black holes with equals masses, $M_{1}=M_{2}=1$, initial positions given by $\boldsymbol{x}^{-}=(-1,0,0)$ and $\boldsymbol{x}^{+}=(+1,0,0)$, zero individual spins and zero linear momenta.
The three-dimensional computational domain is given by $\Omega=[-15;15]^{3}$ covered by a $80\times 80\times 80$ uniform grid, while outgoing boundary conditions are imposed at the border of the computational domain. As shown by Dumbser et al. (2024), at time $t=0$ we find it useful to filter the lapse as
@@
where $\alpha_{min}=0.01$, $\epsilon=10^{-4}$. Moreover, during the evolution, all the metric terms are smoothed out so as to avoid metric spikes, but rather reaching a smooth maximum value at $\gamma_{max}\sim 25$. With these caveats, we have solved this test using $\kappa_{1}=0.2$, $\kappa_{2}=0.2$, $c=0$, $\mu=0.1$ and with the gamma–driver activated. The results of our calculations are reported in Fig. 14 and Fig. 15, respectively, where the contour plots of $\gamma_{xx}$, as well as the shift vector field, are shown at four representative times. Fig. 14 refers to the calculations performed with the FD-WENO scheme, while Fig. 15 refers to those obtained with the AFD-WENO scheme.
By the time $t\sim 10$, the merger of the two black holes is almost complete, and a single stationary black hole is hence produced. The vector field of the shift $\beta^{i}$ manifests the expected behaviour, being oriented in opposite direction with respect to the merger.
@@ 5 Conclusions
Advances in the sensitivities of gravitational wave interferometers are such that they will make it possible to make very accurate observations. As a result, we will soon need very high order schemes for numerical relativity that can match those accurate observations. While second order (in space derivatives) formulations, such as BSSNOK, are routinely adopted, the numerical technology for seamlessly formulating them at very high orders of spatial and temporal accuracy is lacking. This forces us to look at suitable high order accurate discretizations of first order formulations of the Einstein field equations. Because first order formulations have a hyperbolic PDE structure that is common to other very well-explored PDE systems, higher order solution methods have been very well-developed for this class of problem. Therefore, it is very attractive to look at high accuracy methods for solving first order hyperbolic PDE systems. The well-developed higher order metholologies include DG schemes, finite volume WENO schemes and finite difference WENO schemes. All three of these classes of schemes present natural pathways to high order of accuracy in space. However, the first order formulations for numerical GR have the additional complication that they result in extremely large non-conservative hyperbolic PDE systems. These systems are indeed so large that the memory requirements for DG, and perhaps even finite volume WENO schemes, are quite prohibitive. This forces us to take a hard look at recent advances in finite difference WENO schemes.
@@
We have therefore implemented the FD-WENO and AFD-WENO schemes in a three dimensional code that uses Cartesian coordinates within the standard $3+1$ foliation of spacetime. We have successfully reproduced some of the most relevant standard tests of NR in vacuum, as outlined in Alcubierre et al. (2004), which include a gauge wave, the robust stability test, the Gowdy wave, stationary isolated black holes (with excision) and a simple head-on collision of two puncture black holes (Ansorg et al., 2004). Concerning the gauge wave, we could evolve it on a long time scale with no deviations from the analytic solution, provided the order of the scheme is at least 4. This points once more to the conclusion that high order of accuracy might become absolutely crucial in NR. Concerning single stationary black holes, we could evolve them stably over long timescales, thus confirming the positive results obtained by Dumbser et al. (2024) who also adopted the same FO-Z4 formulation, but with an entirely different numerical scheme. In this test case, just like in Dumbser et al. (2024), we had to augment our numerical scheme with a suitable well-balancing feature, in such a way to preserve perfect stationarity. We recall that, on the contrary, the standard BSSNOK formulation has known problems in performing such a test. Academic as it may seem at first sight, this test is nevertheless very important since it opens the door to a rather accurate numerical study of the normal modes of oscillations of black holes (Berti et al., 2009; Mamani et al., 2022). Finally, concerning binary black holes mergers, we confirm the ability of the FO-Z4 formulation in treating simple set-ups with two moving punctures. However, the lack of the conformal factor, that is typical of this formulation, must be compensated by appropriate filtering of the metric spikes that might be produced during the merger. Further investigations are required to make this approach more efficient while at the same time physically neutral.

### v2

@@ 4.4 Stationary black holes
Finally, the bottom right panels of all figures from Fig. 11 to Fig. 14 monitor the evolution of the constraint violations $H(t)-H(0)$ and $M_{i}(t)-M_{i}(0)$ for the Hamiltonian and the momentum constraints. As it can be seen, the effects of the introduced perturbation decay exponentially and the solution converges back to the exact equilibrium. These results confirm what also found by Dumbser et al. (2024).
(a) $\alpha$, lapse profile at $t=1000$ (slice plot) and vector field of the shift.
(b) profile of the lapse error along $x=y=0$.
(c) profile of the $K_{zz}$ error along $x=y=0$.
(d) profile of the $\gamma_{zz}$ error along $x=y=0$.
(e) Evolution of $L_{2}$-error norms for the ADM-constraints
Figure 11: Schwarzschild black hole: Figs. 11a-11d show the absolute errors of $\alpha$, $K_{zz}$ and $\gamma_{zz}$ for the Schwarzschild black hole ($a=0$), computed with respect to the exact solution at the final time $t=1000$. Fig. 11e shows the violation of the ADM-constraints. A seventh-order accurate FD-WENO scheme has been used.
(a) $\alpha$, lapse profile at $t=1000$ (slice plot) and vector field of the shift.
(b) profile of the lapse error along $x=y=0$.
(c) profile of the $K_{zz}$ error along $x=y=0$.
(d) profile of the $\gamma_{zz}$ error along $x=y=0$.
(e) Evolution of $L_{2}$-error norms for the ADM-constraints
Figure 12: Schwarzschild black hole: Figs. 12a-12d show the absolute errors of $\alpha$, $K_{zz}$ and $\gamma_{zz}$ for the Schwarzschild black hole ($a=0$), computed with respect to the exact solution at the final time $t=1000$. Fig. 12e shows the violation of the ADM-constraints. A sixth-order accurate AFD-WENO scheme has been used.
(a) $\alpha$, lapse profile at $t=1000$ (slice plot) and vector field of the shift.
(b) profile of the lapse error along $x=y=0$.
(c) profile of the $K_{zz}$ error along $x=y=0$.
(d) profile of the $\gamma_{zz}$ error along $x=y=0$.
(e) Evolution of $L_{2}$-error norms for the ADM-constraints
Figure 13: Kerr black hole: Figs. 13a-13d show the absolute errors of $\alpha$, $K_{zz}$ and $\gamma_{zz}$ for the Kerr black hole ($a=0.9$), computed with respect to the exact solution at the final time $t=1000$. Fig. 13e shows the violation of the ADM-constraints. A seventh-order accurate FD-WENO scheme has been used.
(a) $\alpha$, lapse profile at $t=1000$ (slice plot) and vector field of the shift.
(b) profile of the lapse error along $x=y=0$.
(c) profile of the $K_{zz}$ error along $x=y=0$.
(d) profile of the $\gamma_{zz}$ error along $x=y=0$.
(e) Evolution of $L_{2}$-error norms for the ADM-constraints
Figure 14: Kerr black hole: Figs. 14a-14d show the absolute errors of $\alpha$, $K_{zz}$ and $\gamma_{zz}$ for the Kerr black hole ($a=0.9$), computed with respect to the exact solution at the final time $t=1000$. Fig. 14e shows the violation of the ADM-constraints. A sixth-order accurate AFD-WENO scheme has been used.
@@ 4.5 Head on collision of two black holes
The correct simulation of binary mergers (either black holes or neutron stars) is of course the final goal of numerical relativity, which is proving to be an invaluable scientific tool in the preset gravitational wave era (Abbott et al., 2019). In this work, still at the level of an exploratory calculation, before a dedicated analysis is performed, we have considered the head-on collision of two nonrotating black holes, which are modelled as two moving punctures. We compute the initial conditions through the TwoPunctures initial data code of Ansorg et al. (2004), selecting two black holes with equals masses, $M_{1}=M_{2}=1$, initial positions given by $\boldsymbol{x}^{-}=(-1,0,0)$ and $\boldsymbol{x}^{+}=(+1,0,0)$, zero individual spins and zero linear momenta [^1] As an alternative way to obtain the initial conditions, one could also adopt the analytic prescription introduced by Brill & Lindquist (1963). See also Alcubierre et al. (2003)..
The three-dimensional computational domain is given by $\Omega=[-15;15]^{3}$ covered by a $80\times 80\times 80$ uniform grid, while outgoing boundary conditions are imposed at the border of the computational domain. As shown by Dumbser et al. (2024), to avoid division by zero at time $t=0$ we find it useful to filter the lapse as
@@
where $\alpha_{min}=0.01$, $\epsilon=10^{-4}$. Moreover, during the time evolution, all the metric terms are naturally smoothed out by the numerical dissipation present in our Riemann-solver-based WENO finite difference schemes. Note that the use of approximate Riemann solvers means upwinding, which is dissipative, and that WENO is a nonlinear shock capturing feature for high order schemes that has been deliberately designed to deal with discontinuities and singularities within the discrete solution. This avoids metric spikes, reaching a rather smooth maximum value at $\gamma_{max}\sim 25$. With these caveats, we have solved this test using $\kappa_{1}=0.2$, $\kappa_{2}=0.2$, $c=0$, $\mu=0.1$ and with the gamma–driver activated. The results of our calculations are reported in Fig. 15 and Fig. 16, respectively, where the contour plots of $\gamma_{xx}$, as well as the shift vector field, are shown at four representative times. Fig. 15 refers to the calculations performed with the FD-WENO scheme, while Fig. 16 refers to those obtained with the AFD-WENO scheme.
By the time $t\sim 10$ (with units of time being $M_{1}=M_{2}$), the merger of the two black holes is almost complete, and a single black hole is hence produced, remaining stationary on longer times. The vector field of the shift $\beta^{i}$ manifests the expected behaviour, being oriented in opposite direction with respect to the merger.
@@ 5 Conclusions
Advances in the sensitivities of gravitational wave interferometers are such that they will make it possible to make very accurate observations. As a result, we will soon need very high order schemes for numerical relativity that can match those accurate observations. This motivated us to look at suitable high order accurate discretizations of first order formulations of the Einstein field equations. Because first order formulations have a hyperbolic PDE structure that is common to other very well-explored PDE systems, higher order solution methods have been very well-developed for this class of problems. Therefore, it is very attractive to look at high accuracy methods for solving first order hyperbolic PDE systems. The well-developed higher-order methodologies include DG schemes, finite volume WENO schemes, and finite difference WENO schemes. All three of these classes of schemes present natural pathways to high order of accuracy in space. However, the first order formulations for numerical GR have the additional complication that they result in extremely large non-conservative hyperbolic PDE systems. These systems are indeed so large that the memory requirements for DG, and perhaps even finite volume WENO schemes, are quite demanding. This forces us to take a hard look at recent advances in finite difference WENO schemes.
@@
We have therefore implemented the FD-WENO and AFD-WENO schemes in a three dimensional code that uses Cartesian coordinates within the standard $3+1$ foliation of spacetime. We have successfully reproduced some of the most relevant standard tests of NR in vacuum, as outlined in Alcubierre et al. (2004), which include a gauge wave, the robust stability test, the Gowdy wave, stationary isolated black holes (with excision) and a simple head-on collision of two puncture black holes (Ansorg et al., 2004). Concerning the gauge wave, we could evolve it on a long time scale with very small deviations from the analytic solution, provided the order of the scheme is at least 4. This points once more to the conclusion that high order of accuracy might become absolutely crucial in NR. Concerning single stationary black holes (not of puncture type but rather the exact solution of Kerr (1963)), we could evolve them stably over long timescales, thus confirming the positive results obtained by Dumbser et al. (2024) who also adopted the same FO-Z4 formulation, but with an entirely different numerical scheme. In this test case, just like in Dumbser et al. (2024), we had to augment our numerical scheme with a suitable well-balancing feature, in such a way to preserve perfect stationarity. We recall that, on the contrary, the standard BSSNOK formulation has known problems in performing such a test. Academic as it may seem at first sight, this test is nevertheless very important since it opens the door to a rather accurate numerical study of the normal modes of oscillations of black holes (Berti et al., 2009; Mamani et al., 2022). Finally, concerning binary black holes mergers, we confirm the ability of the FO-Z4 formulation in treating simple set-ups with two moving punctures. However, the lack of the conformal factor, that is typical of this formulation, must be compensated by appropriate filtering of the metric spikes that might be produced during the merger. Further investigations are required to make this approach more efficient while at the same time physically neutral.